OpenStudy (anonymous):

What is the area of a circle with a radius of 2 meters? Use twenty-two sevenths for pi. a.three and one-seventh square meters b.six and two-sevenths square meters c.twelve and four-sevenths square meters d.twenty-five and one-seventh square meters

Parth (parthkohli):

\[\rm \pi \times radius \times radius\]

OpenStudy (anonymous):

\[A=\pi r ^{2}\] r=radius A=area

OpenStudy (anonymous):